### 18 Dec 2023 v1

# WiSegRT: Dataset for Site-Specific Indoor Radio Propagation Modeling with 3D Segmentation and Differentiable Ray-Tracing

Abstract
The accurate modeling of indoor radio propagation is crucial for localization, monitoring, and device coordination, yet remains a formidable challenge, due to the complex nature of indoor environments where radio can propagate along hundreds of paths. These paths are resulted from the room layout, furniture, appliances and even small objects like a glass cup, and also are influenced by the object material and surface roughness. Advanced machine learning (ML) techniques have the potential to take such non-linear and hard-to-model factors into consideration, but extensive and fine-grained datasets are urgently required. This paper presents WiSegRT [^1] https://github.com/SunLab-UGA/WiSegRT, an open-source dataset for indoor radio propagation modeling. Generated by a differentiable ray tracer within the segmented 3-dimensional (3D) indoor environments, WiSegRT provides site-specific channel impulse responses for each grid point relative to the given transmitter location. We expect WiSegRT to support a wide-range of applications, such as ML-based channel prediction, accurate indoor localization, radio-based object detection, wireless digital twin, etc.
Index Terms:
Radio Propagation, Semantic Segmentation, Deep Learning Channel Modeling, Indoor Radio Dataset, Digital Twin.
I Introduction
The pursuit of advancements in wireless communication systems has led to the exploration of innovative technologies such as integrated sensing and communication (ISAC) [1], reconfigurable intelligent surfaces (RIS) [2], and massive multiple-input multiple-output (mMIMO) systems [3]. These technologies are not only expected to enhance communication data rates but also to perform additional tasks like sensing and localization. The key to their success lies in the availability of accurate and comprehensive wireless channel information. However, obtaining this information poses a significant challenge in the field of radio propagation modeling.
In general, radio propagation models are categorized into two types: probabilistic and deterministic models. Probabilistic models are grounded in empirical formulas and utilize statistical data to calibrate parameters for specific environments. These models, primarily based on the distance between the transmitter (Tx) and receiver (Rx), are excellent at rapid signal strength estimation. However, they often fall short of accuracy and providing detailed channel characteristics such as the channel impulse response (CIR), angle of arrival (AoA), and angle of departure (AoD).
Deterministic models, in contrast, consider the exact environment features. Some deterministic models are simplistic like the free space or the two-ray ground reflection model. Others employ computational electromagnetic (CEM) [4] that utilize Maxwell’s equations with appropriate boundary conditions, which is suitable for small-scale and near-field analysis. For radio propagation in large-scale environments and the transmitter’s far-field, ray concept from geometrical optics [5] is often employed. In a homogeneous medium, usually air, rays travel in straight lines, carrying energy and obeying the laws of reflection, refraction, and diffraction upon interacting with objects. This ray concept is the basis for implementing the well-known ray tracing method for both graphic rendering and radio simulation. However, both CEM methods and approximated ray tracing are computationally intensive and time-consuming, which limits their feasibility for real-time applications that also require high-quality results.
To address the issue of high computational load associated with CEM and ray tracing, the advancements in machine learning (ML) have led to the development of ML-based radio propagation models that offer remarkable capabilities and fast computation speed [6, 7, 8, 9, 10, 11, 12]. The effectiveness of neural network depends significantly on both the quantity and quality of the data. Although there are many radio propagation datasets for outdoor or indoor, such as[13, 14], the majority merely employ 2D layouts or primitive 3D faces as the environment input and coarse channel characterizations, like the received signal strength indicator (RSSI). This setting may be adequate for outdoor scenarios, but often underperform in capturing the complexity of indoor environments which can be found in later section.
In this paper, we introduce Wireless Segmented Ray Tracing (WiSegRT), a precise indoor radio dataset designed for various ML tasks in channel modeling. The main features and contributions are as follows:
  • WiSegRT comprises 10 (more to come) realistic and high-definition indoor scenes with furniture, appliances, and decorations. Objects within these scenes possess distinct and detailed material segmentation, such as the glass panel, plastic casing, and metal stand of a television.
  • Radio simulation was conducted in Sionna [15], a state-of-the-art, open-source, GPU-accelerated physical layer communication system simulator developed by NVIDIA. All its components are differentiable and can be concatenated into neural networks for end-to-end tasks.
  • The 3D scenes were created using Blender software and support high-quality rendering since our scenes also have realistic rendering textures. Researchers can generate corresponding visual-based datasets to implement vision-based ML methods.
II Motivation and Dataset Overview
In this section, we introduce the motivation that inspired us to construct this dataset, and also provide a brief overview of the dataset structure.
II-A Motivation
[…]
However, existing channel datasets often neglect intricate details of indoor environments, such as potential reflection faces, scattering areas, and diffraction edges. Therefore, the model designed for naive 2D or 3D environment and trained on simple data may fail to grasp these rich features. Furthermore, the precision of material segmentation is crucial, neglecting it can result in unexpected high-reflective paths and introducing unstable factors to the system. For more delicate ML-based wireless applications, like mMIMO and radio sensing, the accuracy of radio propagation modeling is the key. Thus, we built WiSegRT, which aims to provide a virtual reality for radio propagation, this means researchers can not only obtain comprehensive radio propagation data from the elaborately created environment, but also generate corresponding visual data for vision-aided radio systems like in [12].
II-B Dataset Overview
Figure 1: WiSegRT dataset file structure
The structure of the WiSegRT dataset is illustrated in Fig. 1. This dataset comprises ten distinct 3D scenes, each corresponding to a scene directory named as {Scene_01, Scene_02...}. For each scene, we set many reasonable Tx locations, which are recorded in the files, named as {Scene_01_tx, Scene_02_tx...}, under the main directory. We also create directories to store data for each Tx under each scene directory, named as {tx_01,tx_02...}. Rx locations are evenly distributed in the space, represented by 3D coordinates [x,y,z] with 20 cm resolution, i.e., x = {0m,0.2m,0.4m...}, y = {0m,0.2m,0.4m...}, z = {0m,0.2m,0.4m...}. We also provide 3D models of each scene, as Scene_01 contained in Scene_01_model. Each Tx directory is further divided into two sub-directories. The first one, path, contains the .pkl files which consists of the Tx position, Rx position, attenuation, delay, AoA and AoD for each Tx-Rx pair. The second sub-directory, obj, houses the .obj 3D model files of the geometrical paths between each Tx-Rx pair, which can be loaded in renders to visualize as shown in Fig. 2. For instance, in Scene_01, there are 30 transmitter positions and 3,325 receiver positions, resulting in a total of 99,750 communication pairs.
Figure 2: Flow of WiSegRT dataset generation
III Technical Details
In this section, we give more technical details about WiSegRT dataset. The system parameters are outlined in Table I, and the synthetic data generation flow is shown in Fig. 2.
Table I  System parameters of WiSegRT
Parameter Value Transmitter Antenna Pattern Dipole Transmitter Antenna Polarization Vertical Receiver Antenna Pattern Isotropic Receiver Antenna Polarization Cross Radio Frequency 2.4 GHz Maximum Interaction 4 Num. of Ray Launched 100,000
III-A Radio Propagation and Ray Tracing
Similar to how light radiates from a bulb and is perceived by the human eye, the radio are emitted from the Tx, propagate in the free space, interact with the objects, bounce zero or several times and reach the Rx. This process results in the radio propagating along various paths with distinct delay, attenuation and phase shift, collectively known as the multi-path effect. The receiver must effectively manage this effect to reconstruct the transmitted information.
As we build the environment models in Blender and imported the scenario to Sionna with electromagnetic (EM) properties assigned to each object, also the Tx and Rx positions, we can initiate the ray launching and tracing processes. To identify all feasible geometrical paths in the multi-path-rich environments, we set Sionna to omnidirectionally launch 100,000 rays from the Tx, and the maximum number of bounces is 4, meaning that if a ray does not reach the Rx after interacting with objects within 4 times, it will be disregarded.
After being launched from the Tx, the ray will travel along straight lines in free space and obey the laws of geometrical optics when interacts with objects. In Sionna, the ray has three types of interactions, e.g. reflection, diffraction and scattering. Simple EM primer for understanding these interactions can be found from Sionna’s document and textbooks. Generally, these interactions are highly related to the material EM properties (electric permittivity, magnetic permeability and conductivity), the incidence angle and the wave polarization. Surface roughness is another critical factor for scattering. By the Rayleigh criterion, if the height of surface irregularities is less than one-eighth of the wavelength, the surface can be considered smooth enough to neglect the scattering effects. In our scenarios, with a carrier frequency of 2.4 GHz corresponding to a wavelength of 0.125 m, most surfaces are deemed smooth, allowing us to disregard scattering path generation. However, for millimeter-wave and terahertz (THz) frequencies, surface roughness and scattering patterns must be considered and configured in Sionna. Once all relevant factors are determined, Sionna calculates the attenuation of the $N$ paths using TensorFlow, as shown in Eq. (1).
$\alpha_{i}=\sum\limits_{i=1}^{N}\frac{\lambda}{4\pi}C_{Rx}(\theta_{Rx,i},\varphi_{Rx,i})^{H}T_{i}C_{Tx}(\theta_{Tx,i},\varphi_{Tx,i})$ (1)
The attenuation for each path in the radio propagation model is denoted by $\alpha_{i}$, where $i$ represents the index of the path. The antenna patterns of the Tx and Rx are represented by $C_{Tx}$ and $C_{Rx}$, respectively. These patterns are utilized in determining the gains of the antennas based on the AoA and AoD, which are denoted by $(\theta_{Rx,i},\varphi_{Rx,i})$ and $(\theta_{Tx,i},\varphi_{Tx,i})$. The wavelength of the radio signal is represented by $\lambda$, and the transformations that occur between the Tx and Rx antennas are encapsulated within the matrix $T_{i}$. These transformations include changes in direction and amplitude attenuation due to free space propagation and interactions with objects. The delay $\tau_{i}$ for the $i$ th path is computed based on the total path length and the speed of light.
It is important to note that the latest Sionna system does not support modeling of refraction. This limitation arises from the complexities involved in accurately representing the internal structure and inhomogeneous materials of objects. For large-scale outdoor scenes this is generally not a significant concern, because their objects are almost large size buildings with glass curtain walls, dense concrete walls and intricate inner structures, which the radio can not penetrate through or reflect out from their inner faces [12]. In contrast, for indoor environments, the effects of the object’s internal parts may become more pronounced due to smaller object sizes. Further research is needed to develop accurate measurement and modeling for these interactions.
III-B Radio and Channels
As we have the attenuation and delay of each path, we can then transfer them to the CIR as each path corresponding to an impulse, as shown in Eq. (2).
$h(\tau)=\sum\limits_{i=1}^{N}\alpha_{i}\delta(\tau-\tau_{i})$ (2)
This is the CIR of the physical wireless channel, which can only be “seen” by the radio frequency front-end. For wireless systems, the focus is more on the baseband equivalent channel and its impulse response, which can be written as Eq. (3).
$h_{b}(\tau)=\sum\limits_{i=1}^{N}\alpha_{i}e^{-j2\pi f\tau_{i}}\delta(\tau-\tau_{i})$ (3)
Here, the $f$ refers to the carrier frequency. The main differences are the phase shifts on each path, as indicated by the exponential term $e^{-j2\pi f\tau_{i}}$. This means that for a given scenario where the paths, $\alpha_{i}$ and $\tau_{i}$ are determined, the phase shift caused by carrier frequency differences, like the differences among sub-carriers, may largely change the baseband equivalent channel, which is known as the frequency selective fading.
IV Comparison of the WiSegRT with Different Settings
As the WiSegRT dataset has very detailed models of objects and precise assignment of EM material properties, it introduces significant overhead in generating additional data and further tasks due to the large amount of primitives the object models contain. This level of detail contrasts with most previous works, which often use vague and simple models and may only consider assigning material for large, strong reflective ones, such as window and metal refrigerator.
In this section, we focus on the “Scene_01” to compare the original WiSegRT dataset, denoted as “ori” (original) scenario, with two other scenarios generated under different settings. To validate the necessity of fine 3D models, we constructed a empty version of the same room without any furniture or appliances, denoted as “emp” (empty) scenario, which only contains the concrete wall, glass windows and wood floor with corresponding EM properties, similar to [9]. To assess the impact of the fineness of environment, we created a simplified version of the room that includes only the main furniture with less precise 3D model and material assignments, denoted as “sim” (simplified) scenario. We designate the “ori” as the ground truth, which is the closest representation of reality. The other two scenarios, due to their partial omission of information, are expected to deviate more from actual conditions.
Figure 3: Comparison of total received power at different distances
Fig. 3 presents a comparison of the average RSSI in different Tx-Rx distance ranges. The RSSI of each Tx-Rx pair is calculated from the total received power of all paths. It is evident from the figure that while the average RSSI of the three scenarios do not significantly differ on decibel (dB) scale, there is a notable difference in their standard deviation (SD). As the Tx-Rx pairs are grouped by their distance, the SD of RSSI is predominantly attributed to the environmental multi-path effects and the blockage of the line-of-sight. This SD serves as an indicator of how accurately each scenario represents the real-world characteristics of radio propagation.
In the “sim” scenario, we can find a reduction in SD compared to the ground truth, which can be attributed to the absence of objects and less detailed material segmentation. Besides, the average RSSI values on dB scale also differ a little between the “ori” and “sim” scenarios, suggesting that fine-grained 3D objects and their material EM properties have a considerable impact on RSSI distribution. In the “emp” scenario, the degradation is worse. As we moved out all the inner objects, the simulated RSSI values exhibit significantly less SD at distances beyond 2m, which means this kind of “emp” setting fails to accurately reflect the RSSI distribution in real-world.
V Possible Applications
V-A Vision-based Radio Tracing
Ray tracing methods, originally designed for graphic rendering to construct digital images, can also be applied to radio propagation modeling. However, significant differences exist between graphical rendering and radio simulations.
In terms of EM waves, graphic rendering mainly deals with visible light, which spans a wide frequency band (380 – 750 THz) and contains a complex mixture of wavelengths. Despite the vast spectrum utilized by the human eye, our vision has relatively low temporal and spectral resolution, capturing only 24-bit color information (8 bits each for red, green, and blue) and processing images at roughly 24 frames per second. Conversely, radio frequencies operate within a comparatively “narrow” band and are discrete, requiring higher resolution in both frequency and time domains, such as 15 kHz and 50 $\mu s$ in LTE. This contrast highlights that graphical rendering is more concerned with the statistical characteristics of each path, whereas radio simulations prioritize accurate attenuation, delay, and phase shift of each path.
On the “receiver” side, graphic rendering requires superior angular resolution and millions of rays to render a high-resolution image. But for radio, the receiver focuses more on how the wireless channel distorts the transmitted EM wave which can be represented by the CIR. In normal indoor scenarios, the number of geometrical paths, also the number of impulses of the CIR, rarely exceeds 200. The number of main paths, those with the least attenuation, rarely exceeds 20. Due to these differences, there is a clear need to develop radio tracing methodologies that can efficiently identify main paths rather than launching an excessive number of rays and discarded most of them.
Neural networks have demonstrated the ability to mimic the radio propagation pattern in 2D environment, as indicated in [7]. It is plausible that neural networks can also learn to identify main paths in the orthogonal views of 3D environments. As shown in Fig. 4, our preliminary scheme is take the orthogonal images from 6 direction (+x,-x,+y,-y,+z,-z) as the environment representation with the Tx and Rx locations, while the output images are orthogonal images of the path objects from 3 direction (+x,+y,+z), which is sufficient to recover the paths. Since the rendered images contain information about object structure and material, it is possible for neural network to implicitly learn the correlation between environmental visuals and main paths. Nevertheless, this idea is in its infancy, and the results to date are limited by both ML theoretical and implementation factors.
Figure 4: Radio tracing with vision data and ML
V-B Indoor Radio Sensing
As EM wave has been deployed in radars for sensing from 1930s, employing radio for indoor sensing is a natural progression due to the ubiquity of wireless devices. Numerous innovative studies have leveraged radio waves for indoor sensing applications, like monitoring human pose and even vital signs. As we provide the 3D model of the environment, it support further scenario customization, including the addition of human figures, editing their actions, and even varying levels of perspiration as it largely change the EM properties of skin and clothing. By understanding how radio interact with target and subsequently reach the Rx, we can build more physics-informed approaches rather than just end-to-end models from the noisy measurements.
VI Conclusions
In this paper, we introduced the WiSegRT, a precise indoor radio propagation dataset. Compared with the simple settings deployed by previous works to generate dataset, the WiSegRT dataset provides an enhanced representation of the complex characteristics of real-world radio propagation. With the virtual-reality-like scene models, WiSegRT supports various ML-based wireless tasks, such as the vision-based radio tracing and ray-level indoor radio-sensing discussed in the application section. Future works will focus on developing model editing libraries to support digital-twin-styled tasks.